OpenStudy (anonymous):

Graph F(x) = x2 and G(x) = 3x + 1; graph F - G.

OpenStudy (anonymous):

F(x) = x^2 and G(x) = 3x + 1; graph F - G.

OpenStudy (anonymous):

F - G = x2 - 3x - 1 which is a quadratic equation